# Analysis of a mathematical model for malaria using data-driven approach

### v1

@@
Abstract: Malaria is one of the deadliest diseases in the world, every year millions of people become victims of this disease and many even lose their lives. Medical professionals and the government could take accurate measures to protect the people only when the disease dynamics are understood clearly. In this work, we propose a compartmental model to study the dynamics of malaria. We consider the transmission rate dependent on temperature and altitude. We performed the steady state analysis on the proposed model and checked the stability of the disease-free and endemic steady state. An artificial neural network (ANN) is applied to the formulated model to predict the trajectory of all five compartments following the mathematical analysis. Three different neural network architectures namely Artificial neural network (ANN), convolution neural network (CNN), and Recurrent neural network (RNN) are used to estimate these parameters from the trajectory of the data. To understand the severity of a disease, it is essential to calculate the risk associated with the disease. In this work, the risk is calculated using dynamic mode decomposition(DMD) from the trajectory of the infected people.
@@ 1 Introduction
According to the WHO report in the year 2022; of all the nations Africa bore the maximum malaria burden. Of all the cases in the world, 94 % of the cases are from Africa, which consists of 233 million cases, and 95 % of total global deaths which consist of 58000 deaths (see Figure 1). It can be clearly understood from the statistics that malaria can cause a huge amount of death in the African continent and it is therefore essential to develop a model which can help doctors and health professionals to understand the disease dynamics in a much better way.
Differential equations-based compartmental models are generally used to analyze the disease transmission. In this method, the entire population is divided into non-overlapping sequential compartments, and the flow of people from one compartment to another is studied using ordinary differential equations. Some of the most commonly used compartments are SIR( susceptible, infected, recovered), SIRD (susceptible infected, recovered, death), and SIRDV(susceptible, infected, recovered, death, vaccinated). Out of all these the simplest and most used model is the SIR model. Analyzing a population using compartmental models has a lot of advantages, which include reducing the complexity of disease transmission, and providing a framework for predicting future states, these models are extremely flexible and can adapt extra compartments in the future if there is a requirement, compartmental models provide a clear visualization of the system’s dynamics.
One of the earliest models for malaria analysis was built by Ross and it consists of only two compartments one for infected humans and the other one for infected mosquitoes, the parameters that were incorporated were biting rate, recovery rate, and death rate. One of the main limitations of this model is that the parasite’s latent period was not considered which is one of the most important factors for malaria transmission. To understand more about this, model one may refer to [1, 2, 3, 4]. After the Ross model, MacDonald developed the next model, which consists of three compartments with both susceptible and infected compartments for mosquitoes and only the infected compartment for human beings. In this model, even though the latent period of parasites in the mosquito was considered, however, the latent period of parasites inside humans was not considered which became a major limitation of this model, to understand more about this model one may refer [5]. Finally, we have the Anderson model which has four compartments where for both the species the susceptible and infected compartments were considered, to know more about this model one may read [6]. To know more about mathematical models of malaria one may refer to the review article [7].
@@
In Ogueda et al.[8], a variant of PINN (physics-informed neural network) called DINN (disease-informed neural network) has been used as a deep learning model. The compartmental model SIRD was used to understand the dynamics of the model, where the dynamics of people traveling from one city to another were also considered. The main aim of this work was to predict various parameters, such as rate of transmission, mortality, and recovery, for the cities taken into consideration, along with the rate of movement of people between the cities. In Schiassi et al. [9], the theory of physics-informed learning was combined with the theory of functional connections. For three different compartmental models namely SIR, SIER, and SIERS the transmission rate, recovery rate, and reproduction number were predicted. In Ning et al.[10] using PINN and SEIRD compartmental model a one-week trajectory was calculated using the COVID-19 data for Italy and along with that, the parameters such as transmission rate, recovery rate, and death rate were predicted from 20 Feb to 25 June. In Ning et al.[11] a variant of the PINN called Euler iteration augmented physics informed neural networks was implemented along with the SIRD compartment model to predict the various transmission rates such as transmission rate, death rate, and along with that forecast for 3 days,5 days, and the 7-day forecast were also calculated.
@@
In Deng et al[12], an RNN-based LSTM deep learning model was combined with the SIRJD compartmental model was used. The main agenda of the work is to predict the transmission rate of the COVID pandemic in the US, and using the predicted transmission rate, the trajectory of the disease was predicted for the next 35 and 42 days. In Bousquet et al. [13], a neural network based on LSTM was implemented on the SIRD model of the coronavirus, the transmission and death rate was predicted along with that the 10-week trajectory for all the compartments. This analysis was done for France, the United Kingdom, Germany, and South Korea. In Deng et al.[14], DNN (deep neural network) and RNN with LSTM have been considered, and they were applied to the SIRD model of the omicron virus to predict the transmission parameters. and based on those parameters, the next 28-day trajectory was predicted, and this analysis was carried out in various cities in China, such as Shanghai, Taiwan, and Hong Kong.
In Hu et al.[15], a modified version of PINN was used to predict various transmission parameters for the SICRD model of COVID-19 and, using the predicted parameters, the next 90-day trajectory for the C and D compartments were predicted, where before the prediction, the data was pre-processed using the wavelet transform. Along with that transmission parameters were predicted with data of 90 days of various cities in the US. It was also shown that the proposed PINN model described in the work is better than the conventional PINN.
In Bhuju et al.[16] the temperature dependence of the transmission rate has been studied using the SEIR model for humans and the LSEI model for mosquitoes. Various mathematical analyses, such as the stability of disease-free equilibrium, and the existence of endemic equilibrium have been done. Numerical results have been carried out using various temperatures and it was observed that temperature plays a huge role in the transmission rate. In Keno et al .[17] the temperature dependency of the transmission, parameter is studied using the SIR model for humans and the SI model for mosquitoes various mathematical analyses such as the local stability of the equilibrium points and global stability of the equilibrium points were performed and it was also proved that if the basic reproduction number is less than one, then the DFE is locally and globally asymptotically stable. Finally, the influence of the temperature on the transmission was studied and it was concluded that temperature plays a significant role in the transmission of the disease.
@@
In Proctor et al.[18] DMD (dynamic model decomposition) was used to incorporate the effect of control to extract low-order models from high-dimensional, complex systems. In Alla and Kutz.[19] DMD is implemented to reduce the order of a nonlinear dynamical system. In Andreuzzi et al.[20] an extension of DMD is used for forecasting the future states of parametric dynamical systems. In Watson et al.[21], the Bayesian time series model has been implemented along with random forest to predict the number of cases and deaths using the SIRD compartmental model. Along with that, the 21-day forecast has been done on three different cities namely New York, Colorado, and West Virginia. More works on the malaria mathematical model can be found in [22, 23, 24, 25, 26, 27, 28, 29].
Most of the works that have been carried out till now related to understanding the dynamics of malaria transmission consist of purely mathematical modeling. But this work makes use of the methods of deep learning in analyzing the dynamics of malaria transmission. One of the reasons why the deep learning method is used instead of other methods is because deep learning models are built to resemble the human brain, enabling them to capture complex patterns in data, This unique property makes them more suitable for modeling malaria dynamics.
In this work, a simple ANN is used to predict the trajectory of all five compartments of the model. To predict the parameters associated with malaria transmission ANN, CNN, and RNN are used. The main reason why RNN was used is because RNN forecasts the future value of a quantity by storing large amounts of past data. Finally, using DMD, the risk of the disease was calculated and the main reason why DMD is used is because DMD can draw insights from raw data, unlike other deep learning methods. The main entity that controls the transmission of a disease is the transmission rate and thus in this work, the transmission rate of disease is analyzed considering the temperature and height simultaneously.
@@
This work is divided in the following manner. In section 2 the model related to disease transmission has been formulated. In section 3 the problem statement, the methodology, and the results are discussed. Section 4 is devoted to concluding remarks and future directions.
@@ 2 Model formulation
In this section, we formulate our model. Both human and mosquito populations are considered in the model since malaria involves two different species mosquitoes and humans. This methodology analyzes the population of humans and mosquitoes using compartment models. The human compartment is divided into three compartments susceptible, infected, and recovered (S, I, R) and the mosquito compartment is divided into two compartments which are susceptible and infected (S, I). This is a two-species system thus the population of one of the species is affected by the other. More precisely when infected mosquitoes come in contact with susceptible humans it leads to an increase in the population of infected humans, in the same way when infected humans come in contact with susceptible mosquitoes it leads to an increase in the population of infected mosquitoes. Figure  2 explains the interaction between the human and mosquito population.
@@ 2.1 Disease-free steady state analysis
We define the basic reproduction number for the model (29) as
@@
If $R_{0}<1$, then the disease-free steady state is locally stable
@@
Dividing the first three equations $N_{h}$ and the last two equations with $N_{m}$ in (29), we get
@@
Now our task is to linearise the system (2.2) around the disease-free steady state. After linearization, we obtain the following system
@@
For the system to be stable, all the eigenvalues must be negative real parts. It is clear that three of the eigenvalues are negative; we need to check the roots of the quadratic polynomial for the remaining two eigenvalues.
Now by solving the quadratic equation, we get the roots as
@@
If we observe, the second root is again always negative, and thus we need to find out the condition for which the first root is negative and that is
@@ 2.2 Endemic steady state analysis
Before proceeding with the next theorem it is essential to define the following quantities
@@
If $R_{0}>1$ and if $b,c,d,bc-d>0$ then the endemic steady state is locally stable.
@@
From the above equation, we can see that we have two linear factors and a cubic factor.
The cubic factor could be analyzed using the following theorem:
@@
Now by using the above-stated lemma, we can arrive at our required result
@@ 2.3 Temperature and altitude dependence of the transmission rate
The transmission rate of the disease is heavily dependent on the altitude of the place as well as the temperature of the place and thus to consider both of the factors we write.
@@
where $\beta_{0}$ is the transmission constant of the region, $T$ is the temperature of the place, $h$ is the height of the place, $\eta$ and $\xi$ are the constants associated with the region’s temperature and height respectively.
Here $e^{-\frac{(T-25)^{2}}{\eta^{2}}}$ is used to model the temperature and $e^{-\frac{h^{2}}{\xi^{2}}}(1-e^{-\frac{h^{2}}{\xi^{2}}})$ is used to model the height. The main reason for considering the above two functions is the following
Temperature:
Malaria transmission will be very minimal when the temperature is either extremely high or it is extremely low and thus to model this variation the Gaussian function is used and the reason for shifting it by 25 is because the optimum temperature for mosquito’s existence and malaria transmission is $25\,^{\circ}\mathrm{C}$.
Altitude:
Malaria transmission is completely zero when the altitude is zero since there won’t be any mosquitoes in the sea and in the same way when the altitude is extremely high again the transmission is completely zero since there are no mosquitoes in the space and thus to model both of these conditions the negative exponential function is used in this manner.
@@
$\beta_{0}=10$, and for human we have taken $\eta=200,\xi=20000$ and for mosquitoes, $\eta=400,\xi=40000$. The different temperatures which are used are $25^{\circ}$,$30^{\circ}$,$35^{\circ}$,$40^{\circ}$,$45^{\circ}$.
The effect of temperature on the transmission rate while maintaining a height of 75 m can be observed in Figures 3 and 4, while the effect at a height of 100 m is shown in Figures 5 and 6, and at a height of 125 m in Figures 7 and 8.
@@
Thus in this graph, we have plotted the graphs for various temperatures and we can observe that the transmission rate is the maximum for $25^{\circ}$ which is the least of all the temperatures.
We verify if $25^{\circ}$ is the optimal temperature. We consider the following values for temperature $10^{\circ}$,$18^{\circ}$,$25^{\circ}$.
The effect of temperature on the transmission rate while maintaining a height of 75 m can be observed in Figures 9 and 10, while at a height of 100 m, it can be observed in Figures 11 and 12, and at a height of 125 m in Figures 13 and 14.
@@
Here transmission is analyzed by keeping the temperature constant and varying the height. Values of altitude used are $150,170,200,220,~{}\text{and}~{}250$ m.
The effect of height on the transmission rate can be observed under different temperatures: at $28^{\circ}$ in Figures 15 and 16, at $35^{\circ}$ in Figures 17 and 18, and at $42^{\circ}$ in Figures 19 and 20.
@@
Here we observe that the transmission achieves its highest value when the height is 150m, which is the least value among the chosen values.
To make sure 150 m is the threshold value for the altitude. Here values of altitude considered are 80 m and 100 m.
The effect of height on the transmission rate while maintaining a constant temperature of $28^{\circ}$ can be observed in Figures 21 and 22, while at $35^{\circ}$ it can be seen in Figures 23 and 24, and at $42^{\circ}$ in Figures 25 and 26

### v2

@@
Abstract: Malaria is one of the deadliest diseases in the world, every year millions of people become victims of this disease and many even lose their lives. Medical professionals and the government could take accurate measures to protect the people only when the disease dynamics are understood clearly. In this work, we propose a compartmental model to study the dynamics of malaria. We consider the transmission rate dependent on temperature and altitude. We performed the steady state analysis on the proposed model and checked the stability of the disease-free and endemic steady state. Since the dynamics of a system are influenced by the parameters, we use three different architectures of neural network namely ANNs(artificial neural networks), RNNs(recurrent neural networks), and PINNs(physics-informed neural networks) to estimate the parameters of the SIR-SI dynamical system and then using the estimated parameters, trajectories of the compartments are predicted. To understand the severity of a disease, it is essential to calculate the risk associated with the disease. In this work, the risk is calculated using dynamic mode decomposition(DMD) from the trajectory of the infected people.
@@ 1 Introduction
According to the World Health Organization (WHO) report from 2022, Africa bears the highest burden of malaria among all regions. Notably, 94 % of the global malaria cases, totaling 233 million, occur in Africa, accounting for 95 % of the worldwide fatalities, which number approximately 58,000 (see Figure 1). These statistics highlight the significant mortality impact of malaria on the African continent, underscoring the urgent need to develop models that enable healthcare professionals to better understand the dynamics of this disease.
Differential equation-based compartmental models are commonly employed to analyze disease transmission. This approach divides the population into non-overlapping sequential compartments, allowing for the study of the movement of individuals between compartments using ordinary differential equations. Some of the most widely utilized compartments include SIR (Susceptible, Infected, Recovered), SIRD (Susceptible, Infected, Recovered, Death), and SIRDV (Susceptible, Infected, Recovered, Dead, Vaccinated).
One of the earliest mathematical models for malaria transmission was developed by Ross, comprising two compartments: one for infected humans and another for infected mosquitoes. Key parameters incorporated in this model includes the biting rate, recovery rate, and death rate. Notably, Ross’s model did not account for the parasite’s latent period, a critical factor in malaria transmission dynamics. For further details on Ross’s work, readers can refer to [1, 2, 3, 4]. Subsequently, MacDonald advanced the modeling framework by introducing a three-compartment model that included both susceptible and infected compartments for mosquitoes, alongside a single infected compartment for humans. Although this model considered the latent period of parasites in mosquitoes, it overlooked the latent period within humans, which represents a significant limitation. More information on MacDonald’s model can be found in [5]. Finally, the Anderson model expanded this approach to four compartments, incorporating susceptible and infected compartments for both species. For additional insights into this model, one can refer to [6]. For a comprehensive overview of mathematical models of malaria, the review article in [7] is recommended.
@@
In the study by Ogueda et al. [8], a variant of the physics-informed neural network (PINN), termed the disease-informed neural network (DINN), was employed as a deep learning model. The SIRD compartmental model was utilized to analyze disease dynamics, incorporating the movement of individuals between cities. The primary objective of this work was to predict various parameters, including the rates of transmission, mortality, and recovery for the selected cities, as well as the rate of movement of individuals between them. In Schiassi et al. [9] work, the principles of physics-informed learning were integrated with the theory of functional connections. This approach was applied to three different compartmental models SIR, SIER, and SIERS to predict key metrics such as the transmission rate, recovery rate, and reproduction number. Ning et al. [10] utilized PINN alongside the SEIRD compartmental model to calculate a one-week trajectory based on COVID-19 data for Italy. This study also predicted parameters including the transmission rate, recovery rate, and death rate from February 20 to June 25. In another study by Ning et al. [11], a variant of the PINN known as Euler iteration-augmented physics-informed neural networks was implemented in conjunction with the SIRD compartmental model. This approach aimed to predict various transmission rates, including the transmission rate and death rate, while also providing forecasts for 3-day, 5-day, and 7-day intervals.
@@
Deng et al. [12] integrated an RNN-based LSTM deep learning model with the SIRJD compartmental model to predict the transmission rate of COVID-19 in the United States, forecasting disease trajectories for the next 35 and 42 days. Bousquet et al. [13] applied an LSTM-based neural network to the SIRD model, predicting transmission and death rates and generating a 10-week trajectory for France, the UK, Germany, and South Korea. Deng et al. [14] used DNN and LSTM with the SIRD model for the Omicron variant, predicting transmission parameters and a 28-day trajectory in cities like Shanghai and Hong Kong. Deng and Wang [15] proposed a deep learning approach combining DNN and LSTM to estimate parameters and predict infections and deaths over 28 days, achieving 98 % accuracy for infections and 92 % for deaths. For more deep learning and statistical models, we refer to [16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
In the study by Bhuju et al. [26], the temperature dependence of the transmission rate was analyzed using the SEIR model for humans and the LSEI model for mosquitoes. The authors conducted various mathematical analyses, including the stability of disease-free equilibrium and the existence of endemic equilibrium. Numerical simulations across different temperature scenarios revealed that temperature significantly influences the transmission rate.
Keno et al. [27] examined the temperature dependency of the transmission parameter using the SIR model for humans and the SI model for mosquitoes. Their analysis includes assessments of both local and global stability of equilibrium points. The study demonstrated that when the basic reproduction number is less than one, the disease-free equilibrium is both locally and globally asymptotically stable. Additionally, the impact of temperature on transmission dynamics was investigated, reinforcing the conclusion that temperature plays a critical role in disease transmission.
@@
In the study by Proctor et al. [28], Dynamic Mode Decomposition (DMD) was utilized to incorporate control effects and extract low-order models from high-dimensional, complex systems. Alla and Kutz [29] implemented DMD to reduce the order of a nonlinear dynamical system. Similarly, Andreuzzi et al. [30] extended DMD for forecasting future states of parametric dynamical systems. Watson et al. [31] employed a Bayesian time series model in conjunction with random forests to predict the number of cases and deaths using the SIRD compartmental model, conducting a 21-day forecast for three cities: New York, Colorado, and West Virginia. Additional research on mathematical models of malaria can be found in [32, 33, 34, 35, 36, 37, 38, 39].
Most existing studies on malaria transmission dynamics primarily rely on mathematical modeling. In contrast, this work leverages deep learning methods to analyze the dynamics of malaria transmission. One key advantage of using the neural network approach is that these models are designed to emulate the human brain, allowing them to capture complex patterns in data. This capability makes deep learning particularly well-suited for modeling malaria dynamics.
In this study, a simple Artificial Neural Network (ANN) is employed to predict the trajectories of all five compartments of the model. To estimate parameters related to malaria transmission, ANN, Convolutional Neural Networks (CNN), and Recurrent Neural Networks (RNN) are utilized. The use of RNN is particularly advantageous, as it can forecast future values by storing extensive historical data. Additionally, Dynamic Mode Decomposition (DMD) is applied to assess the risk of the disease; DMD is effective in extracting insights from raw data, which distinguishes it from other deep learning methods. The primary factor influencing disease transmission is the transmission rate, which is analyzed in this work while considering both temperature and altitude simultaneously.
@@
This study is organized as follows: Section 2 formulates the model related to disease transmission. Section 3 presents the problem statement, methodology, and results. Finally, Section 4 offers concluding remarks and outlines future directions.
@@ 2 Model formulation
In this section, we formulate our model, which encompasses both human and mosquito populations, as malaria involves interactions between these two species. This methodology employs compartment models to analyze the dynamics of humans and mosquitoes. The human population is divided into three compartments: susceptible (S), infected (I), and recovered (R). In contrast, the mosquito population is divided into two compartments: susceptible (S) and infected (I).
Given the two-species nature of this system, the population of one species is influenced by the other. Specifically, when infected mosquitoes come into contact with susceptible humans, the number of infected humans increases. Similarly, when infected humans interact with susceptible mosquitoes, the population of infected mosquitoes rises. Figure 2 illustrates the interactions between the human and mosquito populations.
@@ 2.1 Disease-free steady state analysis
We define the basic reproduction number for the model (2.1) as
@@
If $R_{0}<1$, the disease-free steady state is locally stable.
@@
Dividing the first three equations $N_{h}$ and the last two equations with $N_{m}$ in (2.1), we get
@@
Now our task is to linearize the system (2.2) around the disease-free steady state. After linearization, we obtain the following system
@@
For the system to be stable, all the eigenvalues must have negative real parts. It is clear that three of the eigenvalues are negative; we need to check the roots of the quadratic polynomial for the remaining two eigenvalues. By solving the quadratic equation, we get the roots as
@@
If we observe, the second root is always negative, and thus we need to find out the condition for which the first root is negative and that is
@@ 2.2 Endemic steady state analysis
Before proceeding with the next theorem, it is essential to define the following quantities
@@
If $R_{0}>1$ and $b,c,d,bc-d>0$, the endemic steady state is locally stable.
@@
From the above equation, we can see that we have two linear factors and a cubic factor. To analyze cubic factor, let us state the following lemma:
@@
Now by using the above-stated lemma, we can arrive at our required result. ∎
2.3 Numerical validation of the stability theorems
In the previous section, we observed that the disease-free steady state is achieved when $R_{0}<1$. Accordingly, we selected $R_{0}$ such that $R_{0}<1$, and as shown in Figure 3, the infected populations of humans and mosquitoes diminish over time.
Figure 3: When reproduction number is less than 1
@@ 2.4 Temperature and altitude dependence of the transmission rate
The transmission rate of the malaria disease is heavily dependent on the altitude and temperature, and thus to consider both of the factors we write
@@
where $\beta_{0}$ is the transmission constant of the region, $T$ is the temperature of the region, $h$ is the altitude, $\eta$ and $\xi$ are the constants associated with the region’s temperature and height respectively.
Here $e^{-\frac{(T-25)^{2}}{\eta^{2}}}$ is used to model the temperature and $e^{-\frac{h^{2}}{\xi^{2}}}(1-e^{-\frac{h^{2}}{\xi^{2}}})$ is used to model the height. Malaria transmission will be very minimal when the temperature is either extremely high or it is extremely low and thus to model this variation, the Gaussian function is used and the reason for shifting it by $25$ is because the optimum temperature for mosquito’s existence and malaria transmission is $25\,^{\circ}\mathrm{C}$. Also, malaria transmission is completely zero when the altitude is zero since there won’t be any mosquitoes in the sea and in the same way when the altitude is extremely high again the transmission is completely zero since there are no mosquitoes in the space and thus to model both of these conditions the negative exponential function is used in this manner.
@@
$\beta_{0}=10$, and for human we have taken $\eta=200,\xi=20000$ and for mosquitoes, $\eta=400,\xi=40000$. The different temperature values which we used are $25^{\circ}$, $30^{\circ}$, $35^{\circ}$, $40^{\circ}$, $45^{\circ}$.
The effect of temperature on the transmission rate while maintaining a height of 75 m can be observed in Figures 4 and 5, while the effect at a height of 100 m is shown in Figures 6 and 7, and at a height of 125 m in Figures 8 and 9.
@@
Thus in this graph, we have plotted the graphs for various temperatures and we can observe that the transmission rate is the maximum for $25^{\circ}$ which is the least of all the temperatures. Next, we verify if $25^{\circ}$ is the optimal temperature. We consider the following values for temperature $10^{\circ}$, $18^{\circ}$, $25^{\circ}$. The effect of temperature on the transmission rate while maintaining a height of 75 m can be observed in Figures 10 and 11, while at a height of 100 m, it can be observed in Figures 12 and 13, and at a height of 125 m in Figures 14 and 15.
@@
Malaria transmission is also analyzed by keeping the temperature constant and varying the height. Values of altitude used are $150,170,200,220,~{}\text{and}~{}250$ m. The effect of height on the transmission rate can be observed under different temperatures at $28^{\circ}$ in Figures 16 and 17, at $35^{\circ}$ in Figures 18 and 19, and at $42^{\circ}$ in Figures 20 and 21.
@@
We also observe that the transmission achieves its highest value when the height is 150 m, the lowest value among the chosen values. To make sure 150 m is the threshold value for the altitude. Here values of altitude considered are 80 m and 100 m. The effect of height on the transmission rate while maintaining a constant temperature of $28^{\circ}$ can be observed in Figures 22 and 23, while at $35^{\circ}$ it can be seen in Figures 24 and 25, and at $42^{\circ}$ in Figures 26 and 27.